# Machine Unlearning Through Fine-Grained Model Parameters Perturbation

### v3

@@ III-B2 Unlearning Process
As illustrated in Fig.2, small perturbations are introduced to the model $M$, followed by training the unlearning model $M_{UL}$ using $D_{RE}$ over multiple epochs to enable unlearning while still maintaining the performance of the remaining data. In this process, we gradually and iteratively adjust $M_{UL}$, nudging it towards a state that minimizes the influence of the unlearning data. This adaptive training strategy ensures that $M_{UL}$ retains its high predictive accuracy for the remaining data, while effectively ‘forgetting’ the data that is removed.
@@
For the purpose of eliminating the influence of unlearning data on the model, retraining from scratch, despite consuming substantial computational resources and time, can achieve the optimal unlearning effect. As far as the strategy for machine unlearning is concerned, we desire that the unlearning model $M_{UL}$ approximates the retraining model $M_{RE}$, thus ensuring that the unlearning is moving in the right direction. We aim for both the retraining model and the unlearning model to display similar performance when subjected to $D_{UL}$. However, it clearly goes against the initial intention to perform retraining while conducting unlearning. Despite this, for models trained on i.i.d datasets, the retrained model and $M$ are quite similar, as a small amount of data is unlikely to have a profound impact on the model. Therefore, when calculating the similarity of output distributions, we use $M$ in place of the retrained model in the third step.
@@
Jensen-Shannon (JS) divergence is used to guide the model fine-tuning of the unlearning process [45]. As a symmetric and always finite measure, JS divergence provides a reliable measure of the similarity or distance between two probability distributions. Suppose $x\in D_{UL}$, $P(w,x)$ and $P(\theta,x)$ are posterior distributions of retraining model $M_{RE}$ and unlearning model $M_{UL}$, respectively. $JS$ divergence is calculated as following:
$JS(P(\theta,x)||P(w,x))=\frac{1}{2}KL\left(P(\theta,x)||\frac{P(\theta,x)+P(w,x)}{2}\right)$ (7)
$+\frac{1}{2}KL\left(P(w,x)||\frac{P(\theta,x)+P(w,x)}{2}\right)$
@@
where $KL(\cdot)$ is the Kullback-Leibler divergence. Since $P(\theta,x)$ and $P(w,x)$ are the output probability distributions of $M$ and $M_{UL}$ for $D_{UL}$, we have $\sum P(\theta,x)=\sum P(w,x)=1$, thus:
$JS(P(\theta,x)||P(w,x))=-log2+\frac{1}{2}\sum P(\theta,x)\log P(\theta,x)$
$+\frac{1}{2}\sum P(w,x)\log P(w,x)-\log(P(\theta,x)+P(w,x))$ (8)
@@
In the context of machine unlearning, the JS divergence between distributions $P(w,x)$ and $P(\theta,x)$ can quantify the influence of $D_{UL}$ on $M_{UL}$. If the JS divergence is small, the influence of $D_{UL}$ on $M_{UL}$ is deemed to be minor. Therefore, during training, our goal should be to minimize the JS divergence between these distributions to ensure the effectiveness of unlearning. Moreover, given that the value of the JS divergence always falls between 0 and 1, we can intuitively understand the degree of similarity between $M_{UL}$ and $M$.
@@
$L=\frac{1}{|D_{RE}|}\cdot\sum{y_{i,j}}\cdot log(p_{i,j})+\lambda\cdot JS(P(\theta,x)||P(w,x))$ (9)
Where $|D_{RE}|$ is the number of $D_{RE}$, $y_{i,j}$ is the true label of the $j-th$ category of the $i-th$ sample, and $p_{i,j}$ is the probability that the model predicts the $i-th$ sample as the $j-th$ category, $\lambda$ is used to balance the weight between the original loss and the JS divergence.
@@
4:     Pick a data sample $x$ in $D$, the gradient of $w$: $\frac{\partial F(x,w)}{\partial\theta_{i,j}}$
5:     for $w_{i}$ in $w$ do
6:        calculate sensitivity of each parameter:$sensitivity_{w_{i}}\leftarrow\|g_{w_{i}}(x)\|$
@@
8:     $w_{j}\leftarrow w_{j}+\epsilon\times w_{j},w_{j}\in\{Top-K(w)\}$
@@
10:     $mask\leftarrow Rand(w.shape)$
@@
12:     $w\leftarrow(1-mask)\times w+mask\times w_{j},j\in\{Random-k(w)\}$
@@
Top-K and Random-k machine unlearning strategies are summarized in Algorithms 1.
@@ III-C Unlearning Degree Quantification
By employing a fine-grained model parameter perturbation strategy for the machine unlearning process, the unlearning model and the source model $M$ essentially exhibit the characteristic of model indistinguishability. For datasets with i.i.d attributes, quantifying the unlearning effect by examining the accuracy on the remaining data $D_{RE}$ and the unlearning data $D_{UL}$ of the unlearning model can be challenging. As Fig.1 shows the accuracy of both $D_{RE}$ and $D_{UL}$ may be maintained at a fairly satisfactory level, the difference between these two items in such experiments is about only 15% which is quite small. And Fig.6 displays the same observation, which will be discussed in detail in Section IV-C3.
To address the difficulty in measuring the effect of unlearning and further quantify the degree of forgetting caused by each unlearning strategy, we consider breaking the i.i.d attribute of $D_{UL}$.
Therefore, we design a SPD-GAN applying slight distributional perturbations to $D_{UL}$. It’s necessary to ensure that the added perturbation is minimal so that the model performance difference is not too large while also ensuring that the i.i.d attribute is indeed broken. Therefore, the objective is that perturbed unlearning data $D_{p}$ is guaranteed to perform well on the source model $M$, while the performance of the perturbed unlearning data $D_{p}$ on the unlearning model $M_{UL}$ differs significantly from that of the remaining data $D_{RE}$ on $M_{UL}$. That is to say:
@@
where $\eta$ is a coefficient to balance the $M(\cdot)$’s classification accuracy and $M_{UL}(\cdot)$’s unlearning effect. As Equation.1 explained, our goal is to explicitly demonstrate the degree of unlearning in terms of performance. Once completing the training of SPD-GAN, the performance difference between $M(\cdot)$ and $M_{UL}(\cdot)$ on perturbed data $D_{p}$ is indicative of unlearning degree. That is:
@@
Here, we use ‘accuracy’ to characterize performance $P(\cdot,\cdot)$. In other words, we use $D_{p}$ as an approximation of $D_{UL}$, to measure the performance difference of $D_{UL}$ that does not conform to the i.i.d attribute on models before and after unlearning, thereby characterizing it as the degree of unlearning. Assuming total classes of data is $C$, the unlearning degree ranges in $[0,1-1/C]$.
@@
0:  Joint discriminators $M(\cdot)$ and $M_{UL}(\cdot)$, Generator $G(\cdot)$, coefficient $\eta$
1:  while Epochs do
2:     $\tilde{D_{UL}}=G(D_{UL})+D_{UL}$
3:     $y_{S}\leftarrow M(\tilde{D_{UL}})$
4:     $y_{U}\leftarrow M_{UL}(\tilde{D_{UL}})$
5:     $loss_{M(\cdot)}\leftarrow cross\_entropy(y_{S},Y)$
6:     $loss_{M_{UL}(\cdot)}\leftarrow cross\_entropy(y_{U},Y)$
7:     Minimize $loss\leftarrow loss_{M(\cdot)}-\eta*loss_{M_{UL}(\cdot)}$
8:  end while
9:  $D_{p}=G(D_{UL})+D_{UL}$
10:  $degree=P(M,D_{p})-P(M_{UL},D_{p})$
@@ III-D Time Complexity Analysis
Considering $M_{UL}$, its loss function can be represented as: $L(\theta)=L(w)+\delta L(w)$ with $\delta L(w)$ representing the change due to data removal. For a Top-K unlearning strategy, the gradient update is expressed as $\delta w_{k}=-\alpha\cdot\frac{\partial\delta L}{\partial w_{k}}$ and $\delta w_{i}=0$ for $w_{i}\notin w_{k}$, yielding a time complexity of $\mathcal{O}(K)$ for gradient recomputation. Similarly, for the Random-k strategy, the complexity is $\mathcal{O}(k\cdot N)$ with $k\in(0,1)$. Hence, strategies based on partial parameters perturbed have significantly reduced epoch time compared to full retraining since both $\mathcal{O}(K)$ and $\mathcal{O}(k\cdot N)$ is smaller than $\mathcal{O}(N)$.
@@
Proof: Firstly, suppose $\theta$ is the parameter of the model after perturbation, and $\theta_{rand}$ is the parameter of random initialization, thus:
$\centering L(\theta;D_{RE})\leq L(\theta_{rand};D_{RE})\@add@centering$ (14)
Where $L(\theta,D_{RE})$ denotes the loss of the $D_{RE}$ under the parameter $\theta$. As indicated by Equation 14, the initial loss starting from $\theta$ is expected to be lower.
Furthermore, when considering the magnitude and direction of the gradient $\bigtriangledown L(\theta;D_{RE})$, since $\theta$ represents an optimized solution, for the majority of normalized directions $d$ (where its magnitude is standardized), we have:
$\centering d^{T}\bigtriangledown L(\theta;D_{RE})\leq d^{T}\bigtriangledown L(\theta_{rand};D_{RE})\@add@centering$ (15)
Starting from $\theta$, the requisite step size might be reduced, implying that the gradient’s direction remains more consistent.
Given that the gradient direction originating from $\theta$ exhibits stability and that the starting loss value is already minimized, the iterative optimization is poised to converge more swiftly. Suppose the complete retraining process requires $e_{1}$ epochs, while Top-K machine unlearning takes $e_{2}$ epochs, Random-k takes $e_{3}$ epochs. Consequently, we deduce that: $e_{1}\times\mathcal{O}(N)>e_{2}\times\mathcal{O}(K)$, here both $e_{1}$ surpasses $e_{2}$ and $\mathcal{O}(N)$ exceeds $\mathcal{O}(K)$. Same for Random-k, $e_{1}\times\mathcal{O}(N)>e_{3}\times\mathcal{O}(k\cdot N)$.
@@
We use Equation 5 to approximate the sensitivity of each parameter. Since $|w|=N$, the time complexity for both the forward pass and back-propagation is linear in the number of parameters, the overall time complexity for evaluating sensitivity remains $\mathcal{O}(N)$. Once the model structure is determined, the parameters with Top-K sensitivity of a specific model should be calculated only once. This means that the time required to compute the top-K sensitive parameters will be significantly less than the time taken to perform the unlearning.
@@ IV-A Baseline Methods
The impact of the first few layers on the model is mainly to extract the features of the data, and the latter few layers use these features to perform tasks. Following the setting of [10], we set $K=\{5,10\}$, use EU-K and CF-K as two of our baseline methods. To track the acceleration of machine unlearning, we also take retraining into consideration.
@@ IV-B Evaluation Metrics
To adequately assess the performance of machine unlearning, our considerations are twofold: the effectiveness of the unlearning process and the model generalization. Therefore, we propose novel metrics, namely the forgetting rate and memory retention rate. At the same time, we use acceleration ratio to show the speedup and similarity to measure model indistinguishability.
@@ IV-C1 Training settings
We conduct experiments on CIFAR-10 [46] datasets using ResNet18 [47], VGG [48], GoogLeNet [49], and DenseNet [50] models. The training dataset is randomly divided into $D_{UL}$ and $D_{RE}$ with $D_{UL}$ comprising 5%, 10%, 15% and 20%. For Top-K, we set $K=45$ for ResNet18, and for the Random-k, $k=5\%$. The coefficient $\lambda$ in Equation.9 is set to 0.1. We use Adam as optimizer, for EU-K and CF-K, we set learning rate 1e-4. The generator architecture in SPD-GAN consists of (two Conv2d)-BatchNorm-Conv2d-BatchNorm layers for the encoder, followed by ConvTranspose2d-BatchNorm-ConvTranspose2d-Conv2d layers for the decoder. The LeakyReLU and ReLU activation functions are used for the encoder and decoder, respectively. Activation function Tanh is used to limit the output within [-1, 1]. The coefficient $\eta$ in Equation.12 is 0.03. All the experiments are conducted on a Nvidia A100 GPU.

### v4

@@ III-B2 Unlearning Process
As illustrated in Fig.2, small perturbations are introduced to the model $M$, followed by training the unlearning model $M_{UL}$ using $D_{RE}$ over multiple epochs to enable unlearning while still maintaining the performance of the remaining data. In this process, we gradually and iteratively adjust $M_{UL}$, nudging it towards a state that minimizes the influence of the unlearning data. This adaptive training strategy ensures that $M_{UL}$ retains its high predictive accuracy on the remaining data, while effectively ‘forgetting’ the data that is removed.
@@
For the purpose of eliminating the influence of unlearning data on the model, retraining from scratch, despite consuming substantial computational resources and time, can achieve the optimal unlearning effect. As far as the strategy for machine unlearning is concerned, we desire that the unlearning model $M_{UL}$ approximates the retraining model $M_{RE}$, thus ensuring that the unlearning is moving in the right direction. We aim for both the retraining model and the unlearning model to display similar performance when subjected to $D_{UL}$.
However, it clearly goes against the initial intention to perform retraining while conducting unlearning. Despite this, for models trained on i.i.d datasets, the retrained model and $M$ are quite similar, as a small amount of data is unlikely to have a profound impact on the model. Therefore, when calculating the similarity of output distributions, we use $M$ in place of the retrained model in the third step.
@@
Jensen-Shannon (JS) divergence is used to guide the model fine-tuning of the unlearning process [48]. As a symmetric and always finite measure, JS divergence provides a reliable metric for assessing the similarity or distance between two probability distributions. Suppose $x\in D_{UL}$, $P(\omega,x)$ and $P(\theta,x)$ are posterior distributions of retraining model $M_{RE}$ and unlearning model $M_{UL}$, respectively. JS divergence is calculated as following:
$JS(P(\theta,x)||P(\omega,x))=\frac{1}{2}KL\left(P(\theta,x)||\frac{P(\theta,x)+P(\omega,x)}{2}\right)$ (8)
$+\frac{1}{2}KL\left(P(\omega,x)||\frac{P(\theta,x)+P(\omega,x)}{2}\right)$
@@
Where $KL(\cdot)$ is the Kullback-Leibler (KL) divergence. Since $P(\theta,x)$ and $P(\omega,x)$ are the output probability distributions of $M$ and $M_{UL}$ for $D_{UL}$, we have $\sum P(\theta,x)=\sum P(\omega,x)=1$, thus:
$JS(P(\theta,x)||P(\omega,x))=-log2+\frac{1}{2}\sum P(\theta,x)\log P(\theta,x)$
$+\frac{1}{2}\sum P(\omega,x)\log P(\omega,x)-\log(P(\theta,x)+P(\omega,x))$ (9)
@@
In the context of machine unlearning, the JS divergence between distributions $P(\omega,x)$ and $P(\theta,x)$ can quantify the influence of $D_{UL}$ on $M_{UL}$. If the JS divergence is small, the influence of $D_{UL}$ on $M_{UL}$ is deemed to be minor. Therefore, during training, the objective is to minimize the JS divergence between these distributions to ensure the effectiveness of unlearning. Furthermore, since the JS divergence value always lies between 0 and 1, it provides an intuitive measure of the similarity between $M_{UL}$ and $M$.
@@
$L=$ $\frac{1}{|D_{RE}|}\cdot\underset{(x,y)\in D_{RE}}{\sum}\sum{y_{i,j}}\cdot log(p_{i,j})$
$+\lambda\cdot\frac{1}{|D_{UL}|}\sum_{(x\in D_{UL})}JS(P(\theta,x)||P(\omega,x))$ (10)
Where $|D_{RE}|$ is the number of $D_{RE}$, $y_{i,j}$ is the true label of the $j-th$ class of the $i-th$ sample, and $p_{i,j}$ is the probability that the model predicts the $i-th$ sample as the $j-th$ class, $\lambda$ is used to balance the weight between the cross-entropy loss and the JS divergence.
@@
4:     Pick a data sample $x$ in $D$, the gradient of $\omega$: $\frac{\partial F(x,\omega)}{\partial\theta_{i,j}}$
5:     for $\omega_{i}$ in $\omega$ do
6:        calculate sensitivity of each parameter:$sensitivity_{\omega_{i}}\leftarrow\|g_{\omega_{i}}(x)\|$
@@
8:     $\omega_{j}\leftarrow\omega_{j}+\epsilon\times\omega_{j},\omega_{j}\in\{Top-K(\omega)\}$
@@
10:     $mask\leftarrow Rand(\omega.shape)$
@@
12:     $\omega\leftarrow(1-mask)\times\omega+mask\times\omega_{j},j\in\{Random-k(\omega)\}$
@@
Top-K and Random-k machine unlearning strategies are summarized in Algorithm 1.
@@ III-C Unlearning Degree Quantification
By employing a fine-grained model parameter perturbation strategy for the machine unlearning process, the unlearning model becomes nearly indistinguishable from the source model $M$. Quantifying the unlearning effect by examining the accuracy on the remaining data $D_{RE}$ and the unlearning data $D_{UL}$ of the unlearning model can be challenging. As shown in Fig.1, the accuracy of both $D_{RE}$ and $D_{UL}$ may be maintained at a fairly satisfactory level, the difference between these two items in such experiments is about only 15% which is quite small. Similarly, Fig.6 confirms this observation, which is further elaborated in Section IV-C3.
As mentioned earlier, we believe the reasons can be attributed to the i.i.d property of the data and the generalization ability of the model. To address the difficulty in measuring the effect of unlearning and further quantify the degree of unlearning caused by each unlearning strategy, we consider breaking the i.i.d property of $D_{UL}$.
Therefore, we design a SPD-GAN applying slight distributional perturbations to $D_{UL}$. It’s necessary to ensure that the added perturbation is minimal so that the model performance difference is not too large while also ensuring that the i.i.d property is indeed broken. Therefore, the objective is that perturbed unlearning data $D_{p}$ is guaranteed to perform well on the source model $M$, while the performance of the perturbed unlearning data $D_{p}$ on the unlearning model $M_{UL}$ differs significantly from that of the remaining data $D_{RE}$ on $M_{UL}$. That is to say:
@@
Where $loss_{M(\cdot)}$ and $loss_{M_{UL}(\cdot)}$ represent the losses on $M$ and $M_{UL}$, respectively. And $\eta$ is a coefficient that balances the classification effect of $M(\cdot)$ and the unlearning effect of $M_{UL}(\cdot)$. As explained in Eq.1, our goal is to explicitly demonstrate the degree of unlearning in terms of performance. Once completing the training of SPD-GAN, the performance difference between $M(\cdot)$ and $M_{UL}(\cdot)$ on perturbed data $D_{p}$ is indicative of unlearning degree. That is:
@@
Here, we use accuracy to represent performance $P(\cdot,\cdot)$. Specifically, we use $D_{p}$ as an approximation of $D_{UL}$ to evaluate the performance difference on models before and after unlearning, focusing on the non-i.i.d. property of $D_{UL}$. This difference is then defined as the degree of unlearning. Assuming total classes of data are $C$, the unlearning degree ranges in $[0,1-1/C]$.
@@
0:  Pre-trained joint discriminators $M(\cdot)$ and $M_{UL}(\cdot)$, Generator $G(\cdot)$, coefficient $\eta$
1:  Initialize generator $G(\cdot)$ with random weights
2:  while Epochs do
3:     Sample a batch of unlearning data $D_{UL}$ and corresponding labels $Y$
4:     $D_{p}=G(D_{UL})+D_{UL}$
5:     $y_{S}\leftarrow M(D_{p})$
6:     $y_{U}\leftarrow M_{UL}(D_{p})$
7:     $loss_{M(\cdot)}\leftarrow cross\_entropy(y_{S},Y)$
8:     $loss_{M_{UL}(\cdot)}\leftarrow cross\_entropy(y_{U},Y)$
9:     Minimize $loss\leftarrow loss_{M(\cdot)}-\eta*loss_{M_{UL}(\cdot)}$
10:  end while
11:  $D_{p}=G(D_{UL})+D_{UL}$Generate final perturbed data $D_{p}$ for evaluation}
12:  $degree=P(M,D_{p})-P(M_{UL},D_{p})$
@@ III-D Time Complexity Analysis
Considering $M_{UL}$, its loss function can be represented as: $L(\theta)=L(\omega)+\delta L(\omega)$ with $\delta L(\omega)$ representing the change due to data removal. For the Top-K unlearning strategy, the gradient update is expressed as $\delta\omega_{k}=-\alpha\cdot\partial\delta L/\partial\omega_{k}$ and $\delta\omega_{i}=0$ for $\omega_{i}\notin\omega_{k}$, yielding a time complexity of $\mathcal{O}(K)$ for gradient recomputation. Similarly, for the Random-k strategy, the complexity is $\mathcal{O}(k\cdot N)$ with $k\in(0,1)$. Hence, strategies based on partial parameters perturbed have significantly reduced epoch time compared to full retraining since both $\mathcal{O}(K)$ and $\mathcal{O}(k\cdot N)$ is smaller than $\mathcal{O}(N)$.
@@
Proof: Firstly, suppose $\omega^{\prime}$ is the parameter of the model after perturbation, and $\omega_{rand}$ is the parameter of random initialization, thus:
$\centering L(\omega^{\prime};D_{RE})\leq L(\omega_{rand};D_{RE})\@add@centering$ (15)
Where $L(\omega^{\prime};D_{RE})$ denotes the loss of the $D_{RE}$ under the parameter $\omega^{\prime}$. As indicated by Eq.15, the initial loss starting from $\omega^{\prime}$ is expected to be lower.
Furthermore, when considering the magnitude and direction of the gradient $\bigtriangledown L(\omega^{\prime};D_{RE})$, since $\omega^{\prime}$ represents an optimized solution, for the majority of normalized directions $d$ (where its magnitude is standardized), we have:
$\centering d^{T}\bigtriangledown L(\omega^{\prime};D_{RE})\leq d^{T}\bigtriangledown L(\omega_{rand};D_{RE})\@add@centering$ (16)
Starting from $\omega^{\prime}$, the requisite step size might be reduced, implying that the gradient’s direction remains more consistent.
Given that the gradient direction originating from $\omega^{\prime}$ exhibits stability and that the starting loss value is already minimized, the iterative optimization is poised to converge more swiftly. Suppose the complete retraining process requires $e_{1}$ epochs, while Top-K machine unlearning takes $e_{2}$ epochs, Random-k takes $e_{3}$ epochs. Consequently, we deduce that: $e_{1}\times\mathcal{O}(N)>e_{2}\times\mathcal{O}(K)$, here both $e_{1}$ surpasses $e_{2}$ and $\mathcal{O}(N)$ exceeds $\mathcal{O}(K)$. Same for Random-k, $e_{1}\times\mathcal{O}(N)>e_{3}\times\mathcal{O}(k\cdot N)$.
@@
We use Eq.6 to approximate the sensitivity of each parameter, which is the $L_{2}$ norm of the gradient. For each sample $x$, the time to compute the gradient and the norm is typically linear, i.e., it depends on the number of model parameters and the feature dimension of the sample. As previously noted in Theorem 1, the time for one forward pass and back-propagation is $\mathcal{O}(N)$, hence the overall time complexity for evaluating sensitivity remains $\mathcal{O}(N)$. Once the model structure is determined, the parameters with Top-K sensitivity of a specific model should be calculated only once. This means that the time required to compute the Top-K sensitive parameters will be significantly less than the time taken to perform the unlearning.
@@ IV-A Baseline Methods
The impact of the first few layers on the model is mainly to extract the features of the data, and the latter few layers use these features to perform tasks. Following the setting of [12], we set $K=\{5,10\}$, and use EU-K and CF-K as two of our baseline methods. To track the acceleration of machine unlearning, we also take retraining into consideration.
@@ IV-B Evaluation Metrics
To adequately assess the performance of machine unlearning, we consider two key aspects: the effectiveness of the unlearning process and the model’s generalization. Therefore, we propose novel metrics, namely the forgetting rate and memory retention rate. At the same time, we use acceleration ratio to demonstrate the speedup and similarity to evaluate model indistinguishability.
@@ IV-C1 Training settings
We conduct experiments on CIFAR-10 [49] using ResNet18 [50], VGG [51], GoogLeNet [52], and DenseNet [53] models. The training dataset is randomly divided into $D_{UL}$ and $D_{RE}$ with $D_{UL}$ comprising 5%, 10%, 15% and 20%. For Top-K, we set $K=45$ for ResNet18, and for the Random-k, $k=5\%$. The coefficient $\lambda$ in Eq.10 is set to 0.1. We use Adam as optimizer, for EU-K and CF-K, we set learning rate $41e-$ 4. The generator architecture in SPD-GAN consists of (two Conv2d)-BatchNorm-Conv2d-BatchNorm layers for the encoder, followed by ConvTranspose2d-BatchNorm-ConvTranspose2d-Conv2d layers for the decoder. The LeakyReLU and ReLU activation functions are used for the encoder and decoder, respectively. Activation function Tanh is used to limit the output within [-1, 1]. The coefficient $\eta$ in Eq.13 is 0.03. All the experiments are conducted on a Nvidia A100 GPU.